# FakeNewsGPT4: Advancing Multimodal Fake News Detection through Knowledge-Augmented LVLMs

### v1

@@ Abstract
The massive generation of multimodal fake news exhibits substantial distribution discrepancies, prompting the need for generalized detectors. However, the insulated nature of training within specific domains restricts the capability of classical detectors to obtain open-world facts. In this paper, we propose FakeNewsGPT4, a novel framework that augments Large Vision-Language Models (LVLMs) with forgery-specific knowledge for manipulation reasoning while inheriting extensive world knowledge as complementary. Knowledge augmentation in FakeNewsGPT4 involves acquiring two types of forgery-specific knowledge, i.e., semantic correlation and artifact trace, and merging them into LVLMs. Specifically, we design a multi-level cross-modal reasoning module that establishes interactions across modalities for extracting semantic correlations. Concurrently, a dual-branch fine-grained verification module is presented to comprehend localized details to encode artifact traces. The generated knowledge is translated into refined embeddings compatible with LVLMs. We also incorporate candidate answer heuristics and soft prompts to enhance input informativeness. Extensive experiments on the public benchmark demonstrate that FakeNewsGPT4 achieves superior cross-domain performance compared to previous methods. Code will be available.
FakeNewsGPT4: Advancing Multimodal Fake News Detection through Knowledge-Augmented LVLMs
[…]
1 Introduction
@@
The wide spread of fake news has become an important social issue, posing threats to politics Fisher et al. (2016), finance ElBoghdady (2013), and public health Naeem and Bhatti (2020). Misusing advanced generative models to create misinformation further exacerbates these issues, manifested in the rise of visual deepfakes Zhou et al. (2021) and the proliferation of text fake news Sheng et al. (2021). Furthermore, fake news through multimedia convergence disseminates more expansive information with greater impact to mislead readers. Thus, it is of utmost urgency to develop methods for multimodal fake news detection (MFND).
Figure 1: Illustration of leveraging world knowledge from LVLMs to mitigate the issue of domain shift. (a) Distribution discrepancies of fake news manifested in both manipulated content and realistic context. (b) Profound understanding of unseen fake news aided by the world facts from the LVLMs’ knowledge bases.
Current MFND methods Ying et al. (2023); Chen et al. (2023); Shao et al. (2023a) primarily confine training to given domains (i.e., closed systems), leading to a limited acquisition of information. However, open-world fake news exhibits substantial distribution discrepancies Zhu et al. (2022b), termed domain shift Nan et al. (2021); Zhu et al. (2022c), which is manifested in two aspects in Fig. 1-(a). 1) Diverse manipulated content. Any part of news can be manipulated through creative strategies driven by anonymous intentions. These strategies can selectively distort emotional tendencies, reported incidents, key entities, etc. 2) Diverse realistic context. The real-world context of news differs in regional perspectives and thematic focus. For instance, the common words with distinct styles in BBC are “Prime Minister” and “Queen”, in contrast to “President” and “Obama” in the Washington Post. Hence, it is challenging for existing models in the closed systems to understand and characterize open-world fake news.
Considering the above challenges, we propose to leverage world knowledge from Large Vision-Language Models (LVLMs) Liu et al. (2023); Zhu et al. (2023) to provide complementary information for fake news detectors. This knowledge encompasses a wide array of world facts Bubeck et al. (2023) about public celebrities, social events etc, which enables a comprehensive understanding of unseen fake news. As depicted in Fig 1-(b), the agnostic instance "Donald Trump" can be precisely identified for robust semantic analysis aided by inherent world knowledge from LVLMs.
Despite being proficient in recognizing common instances, LVLMs lack forgery-specific knowledge Zhao et al. (2023), compromising their effectiveness in MFND tasks (see Fig. 2-(a)). Two types of external information are crucial for manipulation reasoning and should be incorporated into the training and inference of vision-language models Zhu et al. (2022a): semantic correlation and artifact trace. On the one hand, manipulated media collectively alter cross-modal correlation, leading to semantic inconsistency between modalities Shao et al. (2023a). On the other hand, manipulated images typically leave suspicious artifact traces Shiohara and Yamasaki (2022).
@@
Figure 2: Illustration of the effect of forgery-knowledge augmentation. (a) Vanilla LVLMs struggle to correctly judge the news veracity. (b) Incorporating external knowledge (i.e., semantic correlation and artifact trace) into LVLMs helps models make accurate predictions.
@@
To enhance cross-domain performance for multimodal fake news detection, in this paper, we present a novel framework, namely FakeNewsGPT4 which augments LVLMs with forgery-specific knowledge. This framework incorporates external information that is beneficial for manipulation reasoning and capitalizes on the world knowledge from LVLMs as complementary. As shown in Fig. 2-(b), we first establish representations of forgery-specific knowledge by cross-modal reasoning and fine-grained verification. Specifically, the multi-level cross-modal reasoning module utilizes cross-modal interactions to focus on regions relevant to semantic correlations. Additionally, the dual-branch fine-grained verification module gathers local spatial information to encode artifact traces through supervised guidance. Subsequently, the generated knowledge is transformed into refined embeddings to align with the language space of LVLMs. We devise MFND instruction-following data for fine-tuning and employ both candidate answer heuristics and soft prompts to unleash the extensive knowledge of language models.
@@
  • We pioneer leveraging world knowledge from large vision-language models (LVLMs) to tackle the domain shift issue in multimodal fake news detection. Building upon capitalizing on the knowledge from LVLMs, we propose a generalized detector, FakeNewsGPT4.
  • FakeNewsGPT4 augments LVLMs with forgery-specific knowledge for manipulation reasoning. We propose cross-modal reasoning and fine-grained verification to extract both semantic correlations and artifact traces. To merge the generated knowledge into LVLMs, we also devise MFND instruction data along with answer heuristics and soft prompts.
@@ 2 Related Work
Figure 3: Architecture of our proposed FakeNewsGPT4. Given a manipulated image-text pair, the multi-level cross-modal reasoning module (a) first extracts cross-modal semantic embeddings and visual patch features. Then, these visual patch features are processed by the dual-branch fine-grained verification module (b) to encode discriminative artifact embeddings. Finally, the semantic and artifact embeddings are incorporated into the forgery-aware vision-language model (c) combined with candidate answer heuristics and soft prompts for deep manipulation reasoning.
Fake News Detection.
@@
Fake news detection works can be categorized into unimodal (image-based and text-based) and multimodal methods. Image-based method Cao et al. (2020) proposes to exploit edited traces to verify the truth of visual content. One group of methods focuses on the spatial domain to mine artifact traces, such as blending Shiohara and Yamasaki (2022), patch consistency Zhao et al. (2021), and reconstruction Liang et al. (2022). The other works transform images into the frequency domain by applying DCT Qian et al. (2020) and phase spectrum Liu et al. (2021b).
Text-based methods primarily delve into various aspects such as sentiment Ghanem et al. (2021), user feedback Min et al. (2022), news environment Sheng et al. (2022), propaganda techniques Huang et al. (2023) and temporal patterns Hu et al. (2023). Recent studies have emphasized the importance of multi-domain Nan et al. (2021); Zhu et al. (2022c) and cross-domain detection methods Nan et al. (2022); Lin et al. (2022) in alleviating the domain shift issue.
@@
In contrast to unimodal methods, multimodal methods adhere to incorporate cross-modal features to extract semantic representations by graph convolution Wang et al. (2020), enhanced entity Qi et al. (2021), co-attention layers Wu et al. (2021), contextual attention Qian et al. (2021), ambiguity learning Chen et al. (2022), multi-view Ying et al. (2023), and causal reasoning Chen et al. (2023). A recent work Shao et al. (2023a) presents HAMMER, a powerful model that combines contrastive learning and cross-modal aggregation. However, all these methods are typically trained and deployed within closed systems, overlooking the potential benefits of accessing world knowledge.
Large Vision-Language Models.
@@
Large language models (LLMs) Touvron et al. (2023); Chiang et al. (2023) have showcased remarkable performance on various linguistic tasks. More recently, researchers are exploring extending the capability of LLMs to perceive visual signals. LLaVA Liu et al. (2023) and Mini-GPT4 Zhu et al. (2023) first facilitate image-text feature alignment followed by visual instruction tuning. PandaGPT Su et al. (2023) employs a simple linear layer as a bridge between ImageBind Girdhar et al. (2023) and the Vicuna model Chiang et al. (2023), allowing for the multimodal input.
Knowledge Augmented Language Models.
@@
Utilizing external knowledge to augment language models (LMs) has emerged as a promising solution in knowledge-intensive tasks Kang et al. (2023). One line of works is retrieval-augmented LMs which retrieve relevant passages and incorporate them into LMs using chunked cross-attention Borgeaud et al. (2022), dual instruction tuning Lin et al. (2024) and self-reflection Asai et al. (2024). Nevertheless, the potential of LVLMs augmented with forgery-specific knowledge in multimodal fake news detection remains unexplored.
3 Proposed Method
@@
In this section, we present a unified framework named FakeNewsGPT4 which incorporates forgery-specific knowledge into LVLMs for manipulation reasoning. We first introduce the overall framework architecture. Then, we elaborate on our proposed multi-level cross-modal reasoning module, dual-branch fine-grained verification module, and forgery-aware vision-language model. Finally, we detail the loss function for training.
3.1 The Overall Framework
@@
The comprehensive architecture of FakeNewsGPT4 is illustrated in Fig. 3. Given a manipulated image-text pair $\left(I,T\right)$, we design two separate modules to extract semantic correlations and artifact traces as forgery-specific knowledge representations. Specifically, the multi-level cross-modal reasoning module utilizes dual-branch cross-attention mechanisms to guide cross-modal interactions, facilitating the encoding of semantic embeddings. Concurrently, the dual-branch fine-grained verification module collects local spatial information to establish artifact embeddings through supervised verification. Then, the forgery-aware vision-language model leverages both the forgery-specific knowledge and inherent world knowledge for deep manipulation reasoning.
3.1.1 Multi-level Cross-modal Reasoning
@@
The input image and text are encoded first by the frozen pre-trained encoders, denoted as $E_{v}$ for the image and $E_{t}$ for the text. To obtain both low-level elements and high-level semantic cues, we partition the image encoder into multiple layers, enabling the integration of intermediate patch-level features. As shown in Fig. 3-(a), the image features $f_{v}$ and text features $f_{t}$ are separately represented as:
$f_{v}=E_{v}\left(I\right),\,\,f_{t}=E_{t}\left(T\right).$ (1)
@@
Both features contain information of object instances within a single modality but lack prior insight into objects referenced by the other modality. This absence of complementary information between modalities may hinder cross-modal semantic reasoning. To this end, we devise the dual-branch cross-attention to guide the interaction between visual and textual features, enabling the extraction of semantic correlations. Attention function Vaswani et al. (2017) is performed on normalized query ($Q$), key ($K$), and value ($V$) features as:
$\begin{split}\text{Attention}(Q,K,V)=\text{Softmax}(\frac{K^{T}Q}{\sqrt{D}})V.\end{split}$ (2)
@@
$u_{v}=\text{Attention}\left(f_{v},f_{t},f_{t}\right),$ (3)
$u_{t}=\text{Attention}\left(f_{t},f_{v},f_{v}\right),$ (4)
@@
Based on the cross-modal interaction described above, the [CLS] tokens $u_{v}^{\rm cls}$ and $u_{t}^{\rm cls}$ can naturally focus on the inter-modal semantic correlations. We concatenate these two [CLS] tokens $\{u_{v}^{\rm cls},u_{t}^{\rm cls}\}$ as a joint representation of semantic embedding. Then we use a learnable linear layer to connect this generated knowledge into the word embedding space of LVLMs, facilitating profound semantic reasoning aided by world factual knowledge.
3.1.2 Dual-branch Fine-grained Verification

### v2

@@ Abstract.
The massive generation of multimodal fake news involving both text and images exhibits substantial distribution discrepancies, prompting the need for generalized detectors. However, the insulated nature of training restricts the capability of classical detectors to obtain open-world facts. While Large Vision-Language Models (LVLMs) have encoded rich world knowledge, they are not inherently tailored for combating fake news and struggle to comprehend local forgery details. In this paper, we propose FKA-Owl, a novel framework that leverages forgery-specific knowledge to augment LVLMs, enabling them to reason about manipulations effectively. The augmented forgery-specific knowledge includes semantic correlation between text and images, and artifact trace in image manipulation. To inject these two kinds of knowledge into the LVLM, we design two specialized modules to establish their representations, respectively. The encoded knowledge embeddings are then incorporated into LVLMs. Extensive experiments on the public benchmark demonstrate that FKA-Owl achieves superior cross-domain performance compared to previous methods. Code is publicly available at https://liuxuannan.github.io/FKA_Owl.github.io/.
Multimodal Fake News Detection, Large Vision-Language Model, Knowledge Augmentation
conference: the 32nd ACM International Conference on Multimedia; October 28 - November 1, 2024; Melbourne, Australia.journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681089isbn: 979-8-4007-0686-8/24/10ccs: Security and privacy Social aspects of security and privacyccs: Information systems Multimedia information systems
1. Introduction
@@
The wide spread of fake news has become an important social issue, posing threats to politics (Fisher et al., 2016), finance (ElBoghdady, 2013), and public health (Naeem and Bhatti, 2020). Misusing advanced generative models (Ho et al., 2020; Cui et al., 2023; Li et al., 2023; Cui et al., 2024a, b) to create misinformation further exacerbates these issues, manifested in the rise of both text fake news (Sheng et al., 2021) and also visual deepfakes (Zhou et al., 2021). Furthermore, multimodal forgery media through convergence disseminates more expansive information with greater impact to mislead readers. Detecting such fake news poses a unique challenge due to the existence of manipulations in both image and text modalities.
@@
Figure 1. Illustration of the effect of forgery-knowledge augmentation. (a) An example of a manipulated image-text pair in which Trump’s face is swapped with another person and the positive words “accept an award” is replaced with the negative “lost an argument”. (b) Existing LVLMs struggle to correctly judge the news veracity. (c) Incorporating forgery-specific knowledge (i.e., semantic correlation and artifact trace) into LVLM helps the model make accurate predictions.
Faced with this challenge, current multimodal fake news detection (MFND) methods (Shao et al., 2023a; Ying et al., 2023; Qian et al., 2021) primarily focus on incorporating cross-modal features. While some progress has been made, the acquisition of broad world information remains challenging due to the confinement of training to given domains (i.e., closed systems). However, open-world fake news exhibits substantial distribution discrepancies (Zhu et al., 2022a), termed domain shift (Nan et al., 2021; Zhu et al., 2022b), which is manifested in abundant forgery methods and diverse real-world context. The existence of domain shift increases the difficulty of understanding and characterizing open-world fake news in MFND tasks.
To address this problem, we propose to leverage Large Vision-Language Models (LVLMs) (Liu et al., 2023; Zhu et al., 2023) which possess rich world knowledge for fake news detectors. This knowledge encompasses a wide array of world facts (Bubeck et al., 2023) about public celebrities, social events etc, which enables a comprehensive understanding of agnostic fake news. However, despite their proficiency in recognizing common instances, the performance of applying off-the-shelf LVLMs to the MFND task is not always satisfying. On the one hand, since LVLMs are not inherently tailored for MFND tasks, they are still challenging to understand and discover the subtle cross-modal differences (Qi et al., 2024; Liu et al., 2024a). For example, in detecting manipulated image-text pair in Fig. 1-(a), the model must discern the sentiment tendencies of both image and text, which are typically not present in LVLMs’ training data. On the other hand, LVLMs lack sensitivity to localized spatial details (Yuan et al., 2024). As shown in Fig. 1-(a), when swapping Trump’s face with another person, there exists intrinsic image discrepancies between edited regions and pristine backgrounds, which is hard to perceive with existing LVLMs. Therefore, it is important to augment LVLMs with external information, referred to as forgery-specific knowledge, which is absent from model parameters yet useful for manipulation reasoning.
We identified two types of forgery-specific knowledge essential for manipulation reasoning (Zhu et al., 2022c): semantic correlation and artifact trace. Firstly, manipulated media often disrupt the natural coherence between different modalities, resulting in semantic discrepancies (Shao et al., 2023a). Secondly, alterations in images usually produce distinctive artifact traces, such as irregular blending boundaries and inconsistencies in color sources, among others (Shiohara and Yamasaki, 2022; Xu et al., 2021). As a result, it would be beneficial to incorporate these two types of knowledge into the training and inference of large vision-language models for multimodal fake news detection.
@@
In this paper, we present a novel framework, namely FKA-Owl which augments LVLMs with forgery-specific knowledge to enhance cross-domain performance for multimodal fake news detection. This framework leverages the rich world knowledge of LVLMs, supplementing it with domain-specific knowledge crucial for identifying multimodal fake news. As shown in Fig. 1-(c), to achieve that, we establish representations of the aforementioned two kinds of forgery-specific knowledge with two specialized modules. The cross-modal reasoning module applies dual cross-attention mechanisms to integrate visual and textual information from frozen encoders, aiming to identify semantic inconsistencies. Meanwhile, the visual-artifact localization module focuses on detecting precise visual artifacts at multiple levels of detail, using sparse bounding boxes and detailed mask regions to trace artifacts. Subsequently, the encoded knowledge representation embeddings are mapped to the language space of LVLMs with projectors. We devise MFND instruction-following data for fine-tuning and employ both candidate answer heuristics and soft prompts to unleash the extensive knowledge of language models.
@@
  • We pioneer leveraging rich world knowledge from large vision-language models (LVLMs) and enhancing them with forgery-specific knowledge, to tackle the domain shift issue in multimodal fake news detection. Our proposed method, FKA-Owl, serves as a general detector to bridge the gap.
  • FKA-Owl augments LVLMs with forgery-specific knowledge for manipulation reasoning. We propose two lightweight modules: the cross-modal reasoning module and the visual-artifact localization module to extract semantic correlations and artifact traces, respectively.
@@ 2. Related Work
Figure 2. Architecture of our proposed FKA-Owl, which is built upon the off-the-shelf LVLM consisting of an image encoder and the LLM. Given a manipulated image-text pair, the cross-modal reasoning module (a) first extracts cross-modal semantic embeddings and visual patch features. Then, these visual patch features are processed by the visual-artifact localization module (b) to encode precise artifact embeddings. Finally, the semantic and artifact embeddings are incorporated into the forgery-aware vision-language model (c) combined with image features and the human prompt for deep manipulation reasoning.
2.1. Fake News Detection
@@
Fake news detection works can be categorized into unimodal (image-based and text-based) and multimodal methods. Image-based method (Cao et al., 2020; Yang et al., 2022) proposes to exploit edited traces to verify the truth of visual content. One group of CNN-based methods focus on the spatial domain to capture artifact traces, such as blending (Li et al., 2020a), multiple instance learning (Li et al., 2020b), patch consistency (Zhao et al., 2021), reconstruction (Huang et al., 2020; Liang et al., 2022), and local mining (Dong et al., 2023). The other works transformed images into the frequency domain by applying DCT (Qian et al., 2020), combining phase spectrum (Liu et al., 2021a), and extracting high-frequency noises (Luo et al., 2021b).
Text-based methods primarily delve into various aspects. Ghanem et al. (Ghanem et al., 2021) proposes to incorporate topic and affective information extracted from text. Some social context-based methods leverage user feedback (Min et al., 2022), news environment (Sheng et al., 2022), propaganda techniques (Huang et al., 2023) and temporal patterns (Hu et al., 2023). Recently, Nan et al. (Nan et al., 2021) and Zhu et al. (Zhu et al., 2022b) all discover the domain shift issue caused by the word frequency and emotion etc, and propose domain gate and domain memory bank to enrich domain information, respectively.
@@
In contrast to unimodal methods, multimodal methods adhere to incorporate cross-modal features to extract semantic representations (Jaiswal et al., 2017). Sabir et al. (Sabir et al., 2018) and Wang et al. (Wang et al., 2020) both propose to combine with the external knowledge base to provide complementary semantics information. Qi et al. (Qi et al., 2021) proposes to extract visual entities to understand the high-level semantics of news. Co-attention network (Wu et al., 2021) and contextual attention network (Qian et al., 2021) are both designed to better fuse textual and visual features. Moreover, Ying et al. (Ying et al., 2023) proposes improved Multi-gate Mixture-of-Expert networks (iMMoE) to refine and fuse features extracted from multiple views. Ambiguity learning (Chen et al., 2022) and causal reasoning (Chen et al., 2023) are separately introduced to address the issue of modal disagreement decisions and spurious correlation in data bias. A recent work (Shao et al., 2023a) presents HAMMER, a powerful model that combines contrastive learning and cross-modal aggregation. However, all these methods are typically trained and deployed within closed systems, overlooking the potential benefits of accessing world knowledge.
2.2. Large Vision-Language Models
@@
Large language models (LLMs) such as GPT-3 (Brown et al., 2020), LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023), have showcased remarkable performance on various linguistic tasks. More recently, researchers are exploring extending the capability of LLMs to perceive visual signals. LLaVA (Liu et al., 2023) and Mini-GPT4 (Zhu et al., 2023) first facilitate image-text feature alignment followed by visual instruction tuning. Visual instruction tuning entails additional training of pre-trained models using curated instruction-formatted datasets to enhance models’ generalization to unseen tasks. PandaGPT (Su et al., 2023) employs a simple linear layer as a bridge between ImageBind (Girdhar et al., 2023) and the Vicuna model, allowing for the multimodal input. The success of LVLMs in the general domain has led to the growth of communities such as medical (Li et al., 2024), video understanding (Jin et al., 2024; Liu et al., 2024b) and image editing (Fu et al., 2024). In this work, we leverage the world knowledge inherent in LVLMs for a better understanding of open-world fake news.
2.3. Knowledge Augmented Language Models
@@
Utilizing external knowledge to augment language models (LMs) has emerged as a promising solution in knowledge-intensive tasks (Kang et al., 2024). One line of works is retrieval-augmented LMs which retrieve relevant passages and incorporate them into LMs. Borgeaud  et al. (Borgeaud et al., 2022) proposes a chunked cross-attention module to incorporate the retrieved text as explicit memory. A lightweight retrieval-augmented dual fine-tuning (Lin et al., 2024) is introduced to retrofit any LLM. Asai et al. (Asai et al., 2024) introduces self-reflection on retrieved passages to enhance LM’s quality and factuality. Nevertheless, the potential of LVLMs augmented with forgery-specific knowledge in multimodal fake news detection remains unexplored.
3. Proposed Method
@@
In this section, we present a unified framework named FKA-Owl which incorporates forgery-specific knowledge into LVLMs for manipulation reasoning. We first introduce the overall framework architecture. Then, we elaborate on our proposed multi-level cross-modal reasoning module, dual-branch visual-artifact localization module, and forgery-aware vision-language model. Finally, we detail the loss function for training.
3.1. The Overall Framework
@@
The comprehensive architecture of FKA-Owl is illustrated in Fig. 2. FKA-Owl consists of an image encoder (ImageBind (Girdhar et al., 2023)), a cross-modal reasoning module, a visual-artifact localization module, and a large language model (Vicuna (Chiang et al., 2023)). Given a manipulated image-text pair $\left(I,T\right)$, we design two lightweight modules to extract semantic correlations and artifact traces as forgery-specific knowledge representations, respectively. Specifically, the cross-modal reasoning module utilizes dual-branch cross-attention mechanisms to guide cross-modal interactions, facilitating the encoding of semantic embeddings. Concurrently, the visual-artifact localization module gathers local spatial information to establish artifact embeddings through supervised localization. Then, the forgery-aware vision-language model leverages both the forgery-specific knowledge and inherent world knowledge for deep manipulation reasoning.
3.2. Multi-level Cross-modal Reasoning
@@
The input image and text are encoded first by the frozen pre-trained encoders, denoted as $E_{v}$ for the image and $E_{t}$ for the text. Both $E_{v}$ and $E_{t}$ originate from the ImageBind, which is aligned with the Vicuna in the off-the-shelf LVLM (Su et al., 2023). To obtain both low-level elements and high-level semantic cues, we partition the image encoder into multiple layers with layer index $l$, enabling the integration of intermediate patch-level features. As shown in Fig. 2-(a), the image features $f_{v}$ and text features $f_{t}$ are separately represented as:
(1) $f_{v}=\sum_{l}{E_{v}^{l}\left(I\right)},\,\,f_{t}=E_{t}\left(T\right).$
@@
Both features contain information of object instances within a single modality but lack prior insight into objects referenced by the other modality. This absence of complementary information between modalities may hinder cross-modal semantic reasoning. To this end, we devise the dual-branch cross-attention to guide the interaction between visual and textual features, enabling the extraction of semantic correlations. Attention function (Vaswani et al., 2017) is performed on normalized query ($Q$), key ($K$), and value ($V$) features as:
(2) $\begin{split}\text{Attention}(Q,K,V)=\text{Softmax}(\frac{K^{T}Q}{\sqrt{D}})V.\end{split}$
@@
(3) $u_{v}=\text{Attention}\left(f_{v},f_{t},f_{t}\right),$
(4) $u_{t}=\text{Attention}\left(f_{t},f_{v},f_{v}\right),$
@@
Based on the cross-modal interaction described above, the [CLS] tokens $u_{v}^{\rm cls}$ and $u_{t}^{\rm cls}$ can naturally focus on the inter-modal semantic correlations. We concatenate these two [CLS] tokens $\{u_{v}^{\rm cls},u_{t}^{\rm cls}\}$ as a joint representation of semantic embedding. Then we use a learnable linear layer to project this generated knowledge into the word embedding space of LVLMs, facilitating profound semantic reasoning aided by world factual knowledge.
3.3. Dual-branch Visual-artifact Localization